OpenStudy (anonymous):

Geometric series. 2+6+18+54+162. What's the value of n?

OpenStudy (anonymous):

3n...........

OpenStudy (anonymous):

am i crct mr Fuzzy

OpenStudy (anonymous):

2 times sum n=0 to infinity 3 to the power of n

OpenStudy (anonymous):

\[T_{n}=ar^n-1\]

OpenStudy (anonymous):

162=2*(3^n-1) 162/2=2/2*(3^n-1) 81=3^n-1 \[3^4=3^n-1\] 4=n-1 n=4+1 =5

OpenStudy (anonymous):

hope u understood